4.The profit function and the average cost function for a product of a company are

p(x) = - 4000 + 194x - 0.149x2 in ringgit and ________ 4000

C(x) = 6 - 0.001x + ________

x

ringgit respectively, when x is the quantity sold in units. Find


(a) the total cost function

(b) the revenue function

(c) the quantity which will maximize the profit

(d) the maximum profit

(e) the price at maximum profit


1
Expert's answer
2021-05-07T10:12:03-0400

(a) The total cost function is:

"C(x) = 6x - 0.001x^2 + 4000."

(b) The revenue function is:

"R(x) = p(x) + C(x) = -4000 + 194x - 0.149x^2 + 6x - 0.001x^2 + 4000 = -0.15x^2 + 200x."

(c) The quantity which will maximize the profit is:

p'(x) = 194 - 0.298x = 0,

x = 194/0.298 = 651 units.

(d) The maximum profit is:

"p(x) = -4000 + 194\u00d7651 - 0.149\u00d7651^2 = 59,147.65."

(e) The price at maximum profit is:

p = R(x)/x = -0.15×651 + 200 = 102.35.